The Jones Company has just completed the third year of a five-year MACRS recovery period for a piece of equipment it originalD. Yes, the cost of taking the order is the lost​ after-tax cash flow of $146,489 from selling the machine.

ANSWER ASAP PLEASE
Answer #1

Book value of Equipment = Purchase cost – Accumulated Depreciation for 3 years

= 299000*(100%-20%-32%-19.20%) = 86,112

b.Sales Price = $179000

Gain on Sale = 179000-86112

= $92,888

Tax on gain = 92,888*35% = $32,510.8

After tax sales proceeds = 179000-32510.8

= $146,489.2

i.e. $146,489

c.D.Yes, the cost is lost after tax cash flows flows of $146,489

i.e. the opportunity cost